# Improving Coverage in Combined Prediction Sets with Weighted p-values

### v2

@@ 1 Introduction
In recent years, machine learning models have achieved remarkable accuracy across a wide range of predictive tasks [1, 2, 3]. Understanding the uncertainty associated with each prediction is essential for decision-making in real-world scenarios, but the black-box nature of many machine learning models hinders their deployment in safety-critical applications such as medical diagnosis [4, 5, 6, 7], industrial control systems [8, 9], and extreme weather forecasting [10, 11, 12]. Conformal prediction [13] emerged as a popular wrapper method around machine learning models because it provides a statistically valid quantification of uncertainty, transforming point predictions to prediction sets with finite-sample coverage guarantees, as long as the test data is exchangeable with the training data.
Complex scenarios with multiple predictions—for example, when there are multiple trials, models, or data sources—naturally produce multiple conformal prediction sets (Figure 1, left). A number of methods have been proposed to aggregate prediction sets. For conformal prediction, where the popular split conformal variant [14] introduces a one-time random split, there are multiple methods of aggregating predictions to reduce the randomness over multiple splits: popular examples include cross-conformal [15], CV+ [16], and jackknife+ [16]. These aggregation methods are all symmetric, in that individual prediction sets contribute equally to the aggregate set. Although comparatively less studied, a more general approach is to aggregate sets asymmetrically—that is, to weight sets based on their prior importance to the overall result [17]. Work in both symmetric and asymmetric aggregation establishes that aggregating individual prediction sets with $1-\alpha$ coverage guarantees results in an overall coverage guarantee of $1-2\alpha$ [18].
@@
  • personalized predictions: “for $X$ location, wind shear is more important than sea temperature”
@@
Our work is based on the observation that if the overall coverage guarantee for asymmetric aggregation reflects the contributions of the individual sets, then we can achieve a tighter guarantee than a constant $1-2\alpha$. Consider, as an extreme, the case where almost all the weight is on one set; then, because the coverage of that set dominates aggregation, the overall coverage guarantee should be closer to the $1-\alpha$ guarantee of the dominant individual set. Accordingly, we propose a method for asymmetric aggregation with data-independent weights where the overall coverage guarantee is based on the distribution of weights across the prediction sets. This approach leverages the results of Vovk and Wang [18] on averaging p-values to achieve tighter guarantees when there exists strong asymmetry in the importance of the prediction sets—i.e. when a single prediction set is significantly more important than the rest—with $1-2\alpha$ as the worst-case guarantee when the contributions of the prediction sets are more equal.
Data-independent weights can encode expert priors on the relative importance of the prediction sets, allowing weights to reflect prior information such as the reliability of the predictors used to generate the sets (Figure 1a). Already, these prior weights generalize standard symmetric aggregation. Yet in practice, many applications demand even more flexibility, and require weights that adapt directly to the data: for example, in the popular mixture-of-experts setting [19], expert weights are learned from inputs. To deal with this general case, we additionally propose a method for aggregating prediction sets with data-dependent weights via a linear transformation on the weighted average of p-values associated with each set. Our method allows us to construct a valid aggregate prediction set while also preserving the proportions of the weights assigned to the individual sets. Since data-dependent weights adjust the contributions of the individual sets based on the observed input, our method also achieves adaptive coverage (Figure 1b). We apply our method to the mixture-of-experts setting, and we demonstrate its effectiveness in experiments with real and synthetic data.
@@ 2 Background
We begin by reviewing preliminaries leading up to the basics of conformal prediction, with a focus on the widely-used split conformal method. To provide context for our method based on p-values, we establish the connection between the more typical quantile presentation of conformal prediction and its original p-value presentation [13].
@@
A p-variable [18] is a random variable $P^{\prime}$ such that
@@ 3 Related work
We categorize existing work on combining conformal results into two approaches: methods for selecting or combining prediction sets, and methods for combining p-values derived from multiple runs of the conformal procedure.
Combining prediction sets
@@
Combining conformal prediction sets was first introduced by Lei et al. [20], who propose combining $K$ split conformal prediction sets by taking their intersection. If each prediction set is constructed at a significance level of $\alpha/K$, then a union bound results in an $\alpha$-level intersection set. Unfortunately, the individual prediction sets (constructed to cover $1-\alpha/K$ of the data) tend to be inefficient—that is, their width or volume is large. Although taking the intersection of these sets reduces the size of the resulting prediction set, the authors demonstrate that, under general conditions, the intersection is wider than the individual prediction sets with probability tending to 1 with an asymptotic number of samples.
@@
Rather than combine prediction sets, Yang and Kuchibhotla [21] propose to select the predictor, from a set of $K$ predictors, that returns the best prediction set. They present two methods: one to select the predictor that gives the most efficient prediction set, but with only approximate validity, and one to select the predictor with the most valid prediction set (i.e. with minimal coverage slack), albeit with a width only close to the minimum. Liang et al. [22] point out that the approximate validity of the first method is due to selection bias, and introduce an alternative approach that both selects the most efficient predictor and uses it to construct a valid prediction set.
@@
A recent line of work explores combining conformal prediction sets by majority vote with a $1-2\alpha$ coverage guarantee [23, 24, 17]. In particular, Gasparin and Ramdas [17] characterize the width of the majority vote set and introduce many extensions, including a weighted majority vote method that incorporates prior information in the weights. Our work on combining conformal results by weighted p-values is closely related to this method. However, our formulation allows us to leverage the results of Vovk and Wang [18] to improve the coverage guarantee beyond the standard $1-2\alpha$, and can also be generalized to data-dependent weights to enable adaptive coverage. (See §C.1 for more detail.)
Combining p-values
@@
A substantial body of research has been devoted to developing methods for combining p-values. Here, we focus specifically on approaches that accommodate arbitrary dependence, with a particular emphasis on their applications to conformal prediction. (For a more comprehensive review, see Balasubramanian et al. [25] and DiCiccio et al. [26].)
@@
An early example of combining p-values with arbitrary dependence is the Bonferroni method, where the minimum of a set of p-values is scaled by the number of p-values. This method has multiple extensions [27, 28], and was first applied to conformal prediction by Lei et al. [20].
@@
Rüschendorf [29] finds that twice the average of p-values is a p-value, a result that was later extended to a more general notion of average by Vovk and Wang [18]. Stutz et al. [30] use the result of Rüschendorf [29] to get a $1-2\alpha$ guarantee for the average of p-values, but also propose a novel transformation to get a p-value average with improved coverage. (For a discussion of how to extend their method to apply to our framework, see §C.2.)
@@
Liu and Xie [31] propose the Cauchy combination test, where the test statistic is a weighted sum of Cauchy-transformed p-values. Wu et al. [32] apply the Cauchy combination test to conformal prediction, and show that their method yields asymptotically exact prediction sets under a weak stability assumption and a joint distribution assumption.
@@ 4 Combining conformal prediction sets by weighted p-value
In this section, we extend the method of combining p-values, as proposed by Vovk and Wang [18], to the setting of conformal prediction. Suppose we have $K$ predictors, denoted by $\widehat{\mu}_{1},\dots,\widehat{\mu}_{K}$, where each predictor $\widehat{\mu}_{k}$ is associated with a nonconformity score function $s_{k}$ and calibration set $S_{k}$. For each predictor, we define a p-value function $\widehat{p}_{k}(x,y)$ that measures how well a candidate label $y$ conforms to the predicted outcome for a given $x$. Specifically, for each $k$, the p-value for a given point is given by
@@
with a guaranteed marginal coverage
@@
Consider assigning a weight $v_{k}$ to each p-value function $\widehat{p}_{k}$. The weighted average p-value function is then
@@
Since the work of Vovk and Wang holds for arbitrarily dependent p-values, this method is robust across a wide range of scenarios. In practice, independent weights allow users to incorporate prior knowledge about the relative quality of different predictors [18, 17]. The weighting can reflect, for example, expert insights on which of $K$ models should be prioritized as being more reliable.
From p-values to prediction sets The work of Vovk and Wang is central to our result, allowing us to generalize and improve upon existing conformal guarantees. Still, despite its broad applicability, their work remains underexplored in conformal literature, and the weight-dependent coverage result has not yet been applied to conformal prediction sets. [^1] To the best of our knowledge, Gasparin and Ramdas [17] is the only work so far to use asymmetric set aggregation, and they derive the standard $1-2\alpha$ guarantee of symmetric aggregation methods. We attribute this oversight to several factors.
Vovk and Wang frame their work in terms of merging functions and p-values, without reference to prediction sets or conformal prediction. As a result, subsequent research has similarly focused on p-values and related test statistics, and has limited connection to conformal literature. Meanwhile, conformal literature typically uses a quantile-based construction of prediction sets, rather than a p-value construction, and work in conformal prediction set aggregation typically operates directly on the sets, rather than working with the associated p-value functions. These trends may contribute to the pattern where p-value results are not always propagated to the wider conformal prediction community. Our result shows that the p-value presentation offers unique benefits to prediction set aggregation, and we hope that our work may encourage renewed interest in the connection between p-values and conformal prediction.
@@ 5 Extending to data-dependent weights
In many practical settings, model weights are determined by the observed data rather than fixed in advance. Such data-dependent weights naturally adapt the influence of each model to the characteristics of the input, making them central to applications like ensemble learning and mixture-of-experts. This adaptivity, however, creates a technical challenge: once the weights depend on the data, they also depend on the associated p-variables, and the theory of Vovk and Wang [18] no longer applies. Nevertheless, we can build on their approach to develop a method that allows us to recover a valid coverage guarantee, even when the weights are data-dependent.
@@
By definition, $m^{*}$ is the minimal scaling factor that makes $P_{\text{all}}$ a valid p-variable, and its value is determined by the CDF $F_{P_{\text{all}}}$. Because $F_{P_{\text{all}}}$ is a population quantity that is not accessible in practice, we cannot evaluate $m^{*}$ directly. Instead, we define a computable proxy $\widehat{m}^{*}$ using an empirical CDF constructed from a designated merging set $S_{\text{merge}}$ [^2] So named because $S_{\text{merge}}$ is used to learn the correction that makes the merging function [18] for the weighted average empirically precise..
@@
Data-dependent weights allow the influence of each model to be adjusted based on how well it performs for a specific data point. This approach enables the construction of aggregated prediction sets that are tailored to the characteristics of the given data, which can be viewed as a form of locally conditional coverage. To be clear, true X-conditional coverage—where the coverage guarantee is conditioned on the current input—is impossible without additional distributional assumptions [33, 34]. However, we demonstrate in §6 that data-dependent weights allow us to create aggregated prediction sets with data-adaptive coverage, which can greatly improve conditional validity in practice.
@@ 6 Application: mixture-of-experts
Mixture-of-experts (MoE) is a machine learning framework designed to combine the predictions of multiple specialized models, called experts [19]. Each expert in an MoE focuses on unique aspects of the problem by learning different representations of the input data. A central component of this framework is the routing network, which determines how to combine the experts’ outputs. Specifically, for an input $x$, the routing network of a traditional (soft) MoE assigns weights $W_{k}(x)$ to each expert output $f_{k}(x)$, producing the final prediction as a weighted sum
@@
Intuitively, the routing weights adapt the contribution of each expert’s p-value to the input so that experts with higher predictive relevance for $X_{n+1}$ have greater influence. This data-adaptive weighting yields coverage guarantees that are locally more robust than split conformal with the full MoE predictor (Figure 2).
@@
We refer to our proposed method of aggregating expert p-value functions with learned weights as weighted aggregation. As a baseline, we compare against split conformal with the full MoE predictor, and we refer to this simply as split conformal for brevity.
To evaluate local validity, we compare against conformal quantile regression (CQR) [35], a widely used locally adaptive method. A strength of our framework is that it complements, rather than competes with, other adaptive methods, allowing weighted aggregation to be layered on top of CQR. Accordingly, we also evaluate a hybrid method that combines weighted aggregation with CQR. We assess performance using marginal coverage, worst-slice (WS) coverage [36], and prediction set size.
@@
We present our main experimental results on the local validity of our method on real-world data in Figure 3. These results focus on regression tasks, where we can compare the standard absolute residual and CQR nonconformity scores; we evaluate our method on classification tasks in §B.2.1. To complement this analysis, which focuses on WS coverage, we also assess local validity from another perspective, by examining coverage disparities across demographic groups in the Communities and Crimes dataset [37] in §B.2.2.
Figure 3: Local validity experiments comparing split conformal to weighted aggregation (WA) using absolute residual scores and CQR scores. Each row corresponds to a dataset, with plots for marginal coverage, WS coverage, and prediction set size from left to right. WA consistently improves WS coverage across datasets, which split conformal undercovers.
In the Appendix, we analyze how various factors impact coverage and prediction set size in a series of ablative studies on synthetic data. We investigate: the effects of merging set size $|S_{\text{merge}}|$B.1.1), shared feature information (§B.1.2), and the different weighted aggregation variants (§B.1.3).
The following experiments are conducted with an MoE where the routing network and experts are all linear models for simplicity. All results are averaged over 200 trials.
@@
Weighted aggregation significantly improves WS coverage, coming close to or meeting the nominal level on all datasets. In addition, the gap between marginal and WS coverage is notably smaller for weighted aggregation than for split conformal, indicating that weighted aggregation provides more uniform coverage, even over the challenging regions of the data. These results show that weighted aggregation is useful in applications where local validity is a priority.

### v3

@@ 1 Introduction
In recent years, machine learning models have achieved remarkable accuracy across a wide range of predictive tasks [khan2022transformers, min2023recent, liang2024foundations]. Understanding the uncertainty associated with each prediction is essential for decision-making in real-world scenarios, but the black-box nature of many of these models hinders their deployment in safety-critical applications such as medical diagnosis [richens2020improving, finlayson2021clinician, chua2023tackling, grote2023uncertainty], industrial control systems [kumar2023machine, lawrence2024machine], and extreme weather forecasting [kashinath2021physics, eyring2024pushing, lai2024machine]. Conformal prediction [vovk2005algorithmic] emerged as a popular wrapper method around machine learning models because it provides a statistically valid quantification of uncertainty. In particular, it transforms point predictions to prediction sets with finite-sample coverage guarantees, as long as the test data is exchangeable with the training data. Complex scenarios with multiple predictions—for example, when there are multiple trials, models, or data sources—naturally produce multiple conformal prediction sets (Figure 1).
A number of methods have been proposed to aggregate prediction sets. For conformal prediction, where the popular split conformal variant [papadopoulos2002inductive] introduces a one-time random split, there are multiple methods of aggregating predictions to reduce the randomness over multiple splits: popular examples include cross-conformal [vovk2015cross], CV+ [barber2021predictive], and jackknife+ [barber2021predictive]. These aggregation methods are all symmetric, in that individual prediction sets contribute equally to the aggregate set. Although comparatively less studied, a more general approach is to aggregate sets asymmetrically—that is, to weight sets based on their prior importance to the overall result [gasparin2024merging, gasparin2024conformal]. Work in both symmetric and asymmetric aggregation establishes that aggregating individual prediction sets with $1-\alpha$ coverage guarantees results in an overall coverage guarantee of $1-2\alpha$ [vovk2020combining].
@@
  • personalized predictions: “for $X$ location and $Y$ month, wind shear is more important than sea temperature”
@@
Our work is based on the observation that if the overall coverage guarantee for asymmetric aggregation reflects the contributions of the individual sets, then we can achieve a tighter guarantee than a constant $1-2\alpha$. Consider, as an extreme, the case where almost all the weight is on one set; then, because the coverage of that set dominates aggregation, the overall coverage guarantee should be closer to the $1-\alpha$ guarantee of the dominant individual set. To this end, we propose a method for asymmetric aggregation with data-independent weights where the overall coverage guarantee is based on the distribution of weights across the prediction sets. This approach leverages the results of vovk2020combining on averaging p-values to achieve tighter guarantees when there exists strong asymmetry in the importance of the prediction sets—i.e. when a single prediction set is significantly more important than the rest—with $1-2\alpha$ as the worst-case guarantee when the contributions of the prediction sets are more equal.
Data-independent weights can encode expert priors on the relative importance of the prediction sets, allowing weights to reflect prior information such as the reliability of the predictors used to generate the sets (Figure 1a). Already, these prior weights generalize standard symmetric aggregation. Yet in practice, many applications demand even more flexibility, and require weights that adapt directly to the data: for example, in the popular mixture-of-experts setting [jacobs1991adaptive], expert weights are learned from inputs. To deal with this general case, we additionally propose a method for aggregating prediction sets with data-dependent weights via a linear transformation on the weighted average of p-values associated with each set. Our method allows us to construct a valid aggregate prediction set while also preserving the proportions of the weights assigned to the individual sets. Since data-dependent weights adjust the contributions of the individual sets based on the observed input, our method also achieves adaptive coverage (Figure 1b). We apply our method to the mixture-of-experts setting, and we demonstrate its effectiveness in experiments with real and synthetic data.
@@ 2 Background
We begin by reviewing preliminaries leading up to the basics of conformal prediction, with a focus on the widely-used split conformal method. To provide context for our method based on p-values, we establish the connection between the more typical quantile presentation of conformal prediction and its original p-value presentation [vovk2005algorithmic].
@@
A p-variable [vovk2020combining] is a random variable $P^{\prime}$ such that
@@ 3 Related work
We categorize existing work on combining conformal results into two approaches: methods for selecting or combining prediction sets, and methods for combining p-values.
@@
Combining prediction sets Combining conformal prediction sets was first introduced by lei2018distribution, who propose combining $K$ split conformal prediction sets by taking their intersection. If each prediction set is constructed at significance level $\alpha/K$, then a union bound yields an overall $\alpha$-level intersection set. However, because each individual prediction set targets the more stringent level $\alpha/K$, they tend to be inefficient—that is, their width or volume is large. Although taking the intersection of these sets reduces the size of the resulting prediction set, the authors show that, under general conditions, this intersection can still be wider than an $\alpha$-level split conformal prediction set, with probability tending to 1 with an asymptotic number of samples.
@@
Rather than combine prediction sets, yang2024selection propose selecting, from a set of $K$ predictors, the predictor that returns the best prediction set. They present two methods: one to select the predictor with the most efficient prediction set, but with only approximate validity, and one to select the predictor with the most valid prediction set (i.e. with minimal coverage slack), albeit with width only close to the minimum. liang2024conformal point out that the approximate validity of the first method is due to selection bias, and introduce an alternative approach that both selects the most efficient predictor and uses it to construct a valid prediction set. Moving beyond choosing a single predictor, luoconformity propose constructing an efficient and valid prediction set by a weighted combination of score functions in the multiclass classification case. In contrast to these methods, which commit to a single prediction set rule globally for all inputs, hegazy2025valid propose a pointwise selection method: for each $X$, they (randomly) choose among candidate prediction sets in a stability-aware way that favors smaller sets, while preserving an appropriately adjusted marginal coverage guarantee.
@@
A recent line of work explores combining conformal prediction sets by majority vote with a $1-2\alpha$ coverage guarantee [cherubin2019majority, solari2022multi, gasparin2024merging, gasparin2024conformal]. gasparin2024merging characterize the width of the majority vote set and introduce many extensions, including a weighted majority vote method that incorporates prior information in the weights. Our work on combining conformal results by weighted p-values is closely related to this method. However, our formulation allows us to leverage the results of vovk2020combining to improve the coverage guarantee beyond the standard $1-2\alpha$, and can also be generalized to data-dependent weights to enable adaptive coverage. (See §C.3 for more detail.)
@@
Combining p-values A substantial body of research has been devoted to developing methods for combining p-values. Here, we focus specifically on approaches that accommodate arbitrary dependence, with a particular emphasis on their applications to conformal prediction. (For a more comprehensive review, see balasubramanian2015conformal and diciccio2020exact.)
@@
An early example of combining p-values with arbitrary dependence is the Bonferroni method, where the minimum of a set of p-values is scaled by the number of p-values. This method has multiple extensions [ruger1978maximale, hommel1983tests], and was first applied to conformal prediction by lei2018distribution.
@@
ruschendorf1982random finds that twice the average of p-values is a p-value, a result that was later extended to a more general notion of average by vovk2020combining. stutz2023conformal use the result of ruschendorf1982random to get a $1-2\alpha$ guarantee for the average of p-values, but also propose a novel transformation to get a p-value average with improved coverage. (For a discussion of how to extend their method to apply to our framework, see §C.4.)
@@
liu2020cauchy propose the Cauchy combination test, where the test statistic is a weighted sum of Cauchy-transformed p-values. wu2023multi apply the Cauchy combination test to conformal prediction, and show that their method yields asymptotically exact prediction sets under a weak stability assumption and a joint distribution assumption.
@@ 4 Combining conformal prediction sets by weighted p-value
In this section, we extend the method of combining p-values, as proposed by vovk2020combining, to the setting of conformal prediction. Suppose we have $K$ predictors, denoted by $\widehat{\mu}_{1},\dots,\widehat{\mu}_{K}$, where each predictor $\widehat{\mu}_{k}$ is associated with a nonconformity score function $s_{k}$ and calibration set $S_{k}$. For each predictor, we define a p-value function $\widehat{p}_{k}(x,y)$ that measures how well a candidate label $y$ conforms to the predicted outcome for a given $x$:
@@
with marginal coverage guarantee
@@
We assign a weight $v_{k}$ to each p-value function $\widehat{p}_{k}$, yielding the weighted average p-value function
@@
Since the work of vovk2020combining holds for arbitrarily dependent p-values, this method is robust across a wide range of scenarios. In practice, independent weights allow users to incorporate prior knowledge about the relative quality of different predictors [vovk2020combining, gasparin2024merging]. The weighting can reflect, for example, expert insights on which of $K$ models should be prioritized as being more reliable.
From p-values to prediction sets
The work of vovk2020combining is central to our result, allowing us to generalize and improve upon existing conformal guarantees. Still, despite its broad applicability, their work remains underexplored in conformal literature, and the weight-dependent coverage result has not yet been applied to conformal prediction sets. [^1] To the best of our knowledge, the works of gasparin2024merging [gasparin2024merging, gasparin2024conformal] are the only ones so far to use asymmetric set aggregation, and they derive the standard $1-2\alpha$ guarantee of symmetric aggregation methods. We attribute this oversight to several factors.
vovk2020combining frame their work in terms of merging functions and p-values, without reference to prediction sets or conformal prediction. As a result, subsequent research has similarly focused on p-values and related test statistics, and has limited connection to conformal literature. Meanwhile, conformal literature typically uses a quantile-based construction of prediction sets, rather than a p-value construction, and work in conformal prediction set aggregation typically operates directly on the sets, rather than working with the associated p-value functions. These trends may contribute to the pattern where p-value results are not always propagated to the wider conformal prediction community. Our result shows that the p-value presentation offers unique benefits to prediction set aggregation, and we hope that our work may encourage renewed interest in the connection between p-values and conformal prediction.
@@ 5 Extending to data-dependent weights
In many practical settings, model weights are determined by the observed data rather than fixed in advance. Such data-dependent weights naturally adapt the influence of each model to the characteristics of the input, making them central to applications like ensemble learning and mixture-of-experts. This adaptivity, however, creates a technical challenge: once the weights depend on the data, they also depend on the associated p-variables, and the theory of vovk2020combining no longer applies. Nevertheless, we can build on their approach to develop a method that allows us to recover a valid coverage guarantee, even when the weights are data-dependent.
@@
By definition, $m^{*}$ is the minimal scaling factor that makes $P_{\text{all}}$ a valid p-variable, and its value is determined by the CDF $F_{P_{\text{all}}}$. Because $F_{P_{\text{all}}}$ is a population quantity that is not accessible in practice, we cannot evaluate $m^{*}$ directly. Instead, we define a computable proxy $\widehat{m}^{*}$ using an empirical CDF constructed from a designated merging set $S_{\text{merge}}$ [^2] So named because $S_{\text{merge}}$ is used to learn the correction that makes the merging function [vovk2020combining] for the weighted average empirically precise..
@@
Data-dependent weights allow the influence of each model to be adjusted based on how well it performs for a specific data point. This approach enables the construction of aggregated prediction sets that are tailored to the characteristics of the given data, which can be viewed as a form of locally conditional coverage. To be clear, true X-conditional coverage—where the coverage guarantee is conditioned on the current input—is impossible without additional distributional assumptions [lei2014distribution, vovk2012conditional]. However, we demonstrate in §6 that data-dependent weights allow us to create aggregated prediction sets with data-adaptive coverage, which can greatly improve conditional validity in practice.
@@ 6 Application: mixture-of-experts
Mixture-of-experts (MoE) is a machine learning framework designed to combine the predictions of multiple specialized models, called experts [jacobs1991adaptive]. Each expert in an MoE focuses on unique aspects of the problem by learning different representations of the input data. A central component of this framework is the routing network, which determines how to combine the experts’ outputs. Specifically, for an input $x$, the routing network of a traditional (soft) MoE assigns weights $W_{k}(x)$ to each expert output $f_{k}(x)$, producing the final prediction as a weighted sum
@@
Intuitively, the routing weights adapt the contribution of each expert’s p-value to the input so that experts with higher predictive relevance for $X_{n+1}$ have greater influence. This pairs naturally with weighted p-value aggregation for two reasons: (1) averaging p-values retains graded evidence from each expert, so that the weight-aggregated p-value faithfully reflects expert disagreement; (2) although data-dependent weights generally break p-value validity, the p-value formulation admits a scaling correction (7) that restores finite-sample validity, even though the routing weights are learned from the data. We expand on both these points in §C.1, and we illustrate the resulting gains in local validity in Figure 2 and in the following experiments.
@@
We refer to our proposed method of aggregating expert p-value functions with learned weights as weighted aggregation. As our main baseline, we compare against split conformal with the full MoE predictor, and we refer to this simply as split conformal for brevity.
To evaluate local validity, we compare against conformal quantile regression (CQR) [romano2019conformalized], a widely used locally adaptive method. A strength of our framework is that it complements, rather than competes with, other adaptive methods, allowing weighted aggregation to be layered on top of CQR. Accordingly, we also evaluate a hybrid method that combines weighted aggregation with CQR. We assess performance using marginal coverage, worst-slice (WS) coverage [romano2020classification], and prediction set size.
@@
For clarity of exposition, we present our main experimental results on the local validity of our method for real-world regression tasks in Figure 3, and defer additional ablation studies and further analyses on local validity to the Appendix.
Figure 3 summarizes our main regression results, with CQR as our main adaptive baseline. We also evaluate our method against additional data-adaptive baselines in §B.2.3 and §C.2, and we present results on classification tasks in §B.2.1.
Our primary metric for evaluating local validity is WS coverage. To complement this, we also examine coverage disparities across demographic groups on the Communities dataset [redmond2002data] in §B.2.2
Figure 3: Local validity experiments comparing split conformal to weighted aggregation using absolute residual scores and CQR scores. Each row corresponds to a dataset, with plots for marginal coverage, WS coverage, and prediction set size from left to right. The weighted aggregation methods consistently improves WS coverage across datasets, in most cases meeting target coverage on the WS slab. Split conformal is less conservative marginally, but undercovers on the WS slab; using CQR scores is not enough to cover the gap.
In Appendix B, we analyze how various factors impact coverage and prediction set size in a series of ablation studies on synthetic data. We investigate: the effects of merging set size $|S_{\text{merge}}|$B.1.1), shared feature information (§B.1.2), and the different weighted aggregation variants (§B.1.3).
The following experiments are conducted with an MoE where the routing network and experts are linear models for simplicity. All results are averaged over 200 trials.
@@
Weighted aggregation significantly improves WS coverage, coming close to or meeting the nominal level on all datasets. In addition, the gap between marginal and WS coverage is notably smaller for weighted aggregation than for split conformal, indicating that weighted aggregation provides more uniform coverage, even over the challenging regions of the data. These results suggest that weighted aggregation is preferable in applications where local validity is a priority.